identify the focus and the directrix of the graph of x=-1/20y^2
Answer by lwsshak3(11628)

identify the focus and the directrix of the graph of
x=-1/20y^2
y^2=-20x
This is an equation of a parabola that opens leftward with vertex at the origin.
Its basic form of equation: y^2=-4px
4p=20
p=5
directrix:x=5
focus(-5,0)